OpenStudy (anonymous):

x/12=55/60 a 5 b 7 c 10 d 11 i think its D 11

OpenStudy (kendricklamar2014):

@Wendy.Ivette11714

OpenStudy (wendy.ivette11714):

Same as the last problem: \[x \times(60)=(55)\times(12)\]

OpenStudy (anonymous):

60x and 660?

OpenStudy (wendy.ivette11714):

Yes that is correct

jabez177 (jabez177):

If you think it is D. 11 then put it in the expression and see if it works. \[\frac{ 11 }{ 12 }=\frac{ 55 }{ 60 }\]

OpenStudy (wendy.ivette11714):

Now we divide both sides by 60: \[\frac{ 60x }{ 60 } = \frac{ 660 }{ 60 }\]

jabez177 (jabez177):

What x 11 will equal 55 and what times 12 will equal 60?

OpenStudy (wendy.ivette11714):

Can you solve that @jayjaysimmer

OpenStudy (anonymous):

660 divided by 60 is 11

OpenStudy (wendy.ivette11714):

Yes the answer is x = 11

OpenStudy (anonymous):

Thank You!

OpenStudy (wendy.ivette11714):

Your Welcome :)